OpenStudy (narissa):

Given the formula: TF=95Tc+32 TF represents Temperature in Fahrenheit TC represents Temperature in Celsius Select a correct formula to solve for the Temperature in Degrees Celsius Question 5 options: TC=(TF−32)⋅95 TC=(TF−32)⋅59 TC=(TF+32)⋅95 TC=(TF+32)⋅59

OpenStudy (narissa):

@Lovelarap

OpenStudy (anonymous):

First take 32 to the other side

OpenStudy (hossamhoussien):

Are these options TRUE? as it should be \[t_{c} = (t _{f}-32)/95\] which not exist in the choices

OpenStudy (narissa):

tc=(tf-32)9/5

OpenStudy (narissa):

@HossamHoussien

OpenStudy (anonymous):

um, i think you mean tc=(tf-32)/95

OpenStudy (narissa):

i dont have any answer choices that look like that

OpenStudy (hossamhoussien):

maybe there is a mistake in the question if u can post the original question here as picture

OpenStudy (narissa):

above ^^^

OpenStudy (hossamhoussien):

No, i mean the original formula TF=95Tc+32

OpenStudy (hossamhoussien):

it's 9/5 and u just wrote 95 that why there was an error. anyway the answer would be \[T _{c}=[T _{f}-32]\times \frac{ 5 }{ 9 }\]

OpenStudy (narissa):

oh sorry thanks :))

OpenStudy (anonymous):

The formula for finding the Celsius is TC = TF/1.8 + 32.0.